# AVG-LLaVA: A Large Multimodal Model with Adaptive Visual Granularity

### v2

@@ 4.1 Settings
Training Datasets
Note that in this work, we mainly focus on investigating the effectiveness of automatic visual granularity selection in reducing the number of visual tokens and improving model performance. Therefore, during the first stage, we also use CC-595K image-text pairs for model training, as implemented in LLaVA-NeXT (Liu et al., 2024b). In the subsequent training stages, we also hope to use the same data as LLaVA-NeXT. However, the real user interaction data used for visual instruction fine-tuning in LLaVA-NeXT are not open-sourced, so we opt to extract 200K samples from ALLaVA (Chen et al., 2024a) dataset as a substitute. Although LLaVA-NeXT replaces TextVQA (Singh et al., 2019) with DocVQA (Mathew et al., 2021) and SynDog-EN (Kim et al., 2022), the TextVQA has already been included in the training data of most existing LMMs. Consequently, we choose to retain it to ensure a fair comparison with other models. [^2] Our data recipe refers to Open-LLaVA-NeXT (Lin & Long, 2024) In total, the visual instruction fine-tuning data we use contains 1M image-text pairs.
Implementation Details
Following Liu et al. (2024b), we use CLIP ViT-L/14 as the visual encoder, Vicuna-7B (Chiang et al., 2023) as the LLM. We set the filtered instruction token number $k$ to 32 and the cross-entropy loss weight $\alpha$ to 0.1. The impact of these two hyperparameters on model performance is discussed in Section 4.6. In our first two stages, we adopt the same training settings as LLaVA-NeXT for pre-training and visual instruction fine-tuning. In the third stage, the learning rates for the visual encoder and other modules are set to $2$$\times$10-5 and $1$$\times$10-5, respectively, with a batch size of 128. In the fourth stage, the learning rate for the visual granularity router is set to $1$$\times$$10^{-3}$, with a batch size of 128. More details of the training process are provided in Appendix A.1.
Evaluations
We evaluate our model on three kinds of benchmarks: (1) general VQA benchmarks: GQA (Hudson & Manning, 2019), SciQA-Img (Lu et al., 2022), and VizWiz (Gurari et al., 2018); (2) text-oriented VQA benchmarks: TextVQA (Singh et al., 2019), ChartQA (Masry et al., 2022), DocVQA (Mathew et al., 2021), and AI2D (Kembhavi et al., 2016); and (3) general multimodal benchmarks: MME (Fu et al., 2023), MMB (Liu et al., 2023b), MMBCN (Liu et al., 2023b), POPE (Li et al., 2023c), and MMMU (Yue et al., 2023).
4.2 Main Results
Table 1: Comparison with LMMs of the same size on general VQA benchmarks and text-oriented VQA benchmarks. The best results are marked in bold, and the second best results are underlined.
@@ 4.2 Main Results
Table 2: Comparison with LMMs of the same size on general multimodal benchmarks. The best results are marked in bold, and the second best results are underlined.
@@
General VQA Benchmarks
The results in Table 1 show that AVG-LLaVA outperforms all standard-resolution LMMs on the General VQA benchmarks and other high-resolution LMMs on VizWiz. Although it does not achieve the best results on GQA and ScienceQA, it is important to note that AVG-LLaVA uses fewer visual tokens compared to other high-resolution models, and this comparison will be detailed in Section 4.3.
Text-oriented VQA Benchmarks
In this category of benchmarks, as illustrated in Table 1, except for TextVQA, AVG-LLaVA outperforms all other comparison models. Back to TextVQA, AVG-LLaVA achieves the second-best result, only trailing behind Mini-Gemini-HD. Notably, Mini-Gemini-HD utilizes more than twice the amount of data during the pretraining and approximately 1.5 times the amount of data during the visual instruction fine-tuning compared to AVG-LLaVA.
General Multimodal Benchmarks
Compared to traditional VQA datasets, this type of benchmark covers a broader range of evaluation aspects, requiring models to possess more complex perception and reasoning capabilities. As summarized in Table 2, AVG-LLaVA surpasses all other models, exhibiting superior overall performance and highlighting its adaptability and effectiveness across various disciplines. Specifically, AVG-LLaVA outperforms the second best model by 9.4 and 6.1 on MME and MMEC, respectively, and by 1.9 and 1.2 on MMB and MMBCN, respectively. Moreover, AVG-LLaVA’s performance on the POPE and MMMU benchmarks demonstrates its ability to reduce hallucinations and perform complex reasoning.
4.3 Computational Efficiency
Table 3: Comparisons of AVG-LLaVA and LLaVA-NeXT in terms of the number of visual tokens and actual inference speed, both of which are tested on 8 V100 GPUs with a batch size of 1. AVG-LLaVA can reduce the number of visual tokens by up to 85.3% and is up to 2.53$\times$ faster than LLaVA-NeXT.
@@ 4.3 Computational Efficiency
Speed $\uparrow$ 1.14$\times$ 1.77$\times$ 1.41$\times$ 1.04$\times$ 0.97$\times$ 2.53$\times$ 1.19$\times$ 1.87$\times$ 1.79$\times$
To validate the effectiveness of dynamic visual granularity selection, we compare AVG-LLaVA with LLaVA-NeXT in terms of visual token number and inference speed across multiple benchmarks. Specifically, for each type of benchmark, we select three benchmarks for comparison, and report the reduction in the number of visual tokens per grid and the actual speedup during inference. As shown in Table 3, except for text-intensive VQA benchmarks that require very fine-grained visual information, such as TextVQA and ChartVQA, AVG-LLaVA significantly reduces the number of visual tokens and improves inference speed across other benchmarks. Particularly, on the AI2D benchmark, AVG-LLaVA achieves better performance than LLaVA-NeXT while using only 14.7% of the visual tokens, and the inference speed increased by 2.53 $\times$. [^3] We also present qualitative results in Appendix A.2 and illustrate the effectiveness of adaptive visual granularity. Notably, even with the addition of two extra modules, there is no significant slowdown in inference speed on the ChartVQA benchmark when using a comparable number of visual tokens. Moreover, AVG-LLaVA only increases the number of parameters by 1.66% compared to LLaVA-NeXT.
@@ 4.4 Routing Visualization
To further understand the differences in the granularity selection of AVG-LLaVA across different benchmarks, we visualize the proportion of visual tokens selected at each granularity level for all benchmarks. Figure 5 shows the visualization results, it is evident that different tasks tend to favor different visual granularity, which is consistent with our expectations. In the case of text-intensive benchmarks like TextVQA, ChartQA, and DocVQA, the model requires fine-grained visual information, so the router predominantly selects the finest visual granularity. On the other hand, for benchmarks involving object-level questions, such as AI2D and MMMU, the model may find it easier to answer correctly by utilizing coarse-grained visual information. Additionally, we observe that the granularities with 72 and 288 visual tokens are rarely selected. However, we find that incorporating these granularities of visual tokens helps the model progressively learn to utilize visual tokens with different granularity and allows it to more fully understand the differences between various visual granularities. Further verifications are detailed in Section 4.5.
@@ 4.5 Ablation Study
Adaptive Visual Granularity vs. Fixed Visual Granularity
@@
Router Granularity Selection vs. Random Granularity Selection
@@
Impact of Router Input
@@
Impact of Granularity Range
@@
Impact of Router Training Methods
We directly train the router using visual instructions fine-tuning, applying the same loss function as in Stage 2. Unlike our original approach where the router is directly supervised by LMM feedback, this variant computes the loss on the LMM and backpropagates the gradient to the router using the Gumbel-Softmax technique (Jang et al., 2017). The results in Table 4(e) show that direct feedback from the LMM allows the router to better distinguish the advantages and disadvantages of different granularities, thereby enabling it to select an appropriate granularity.
Importance of Ranking Granularity
In Table 4(f) and Table 4(g), we remove the cross-entropy loss and ranking loss during the fourth stage, respectively. The results indicate that both types of loss are beneficial to model training and are complementary to each other, between which the ranking loss is more crucial. This underscores the necessity to train the router by ranking granularity to align LMM feedback.
4.6 Hyperparameter Analysis
@@
5 Conclusion
In this work, we propose AVG-LLaVA, an LMM that can adaptively select appropriate visual granularity based on input image and instruction. AVG-LLaVA builds upon LLaVA-NeXT by introducing a visual granularity scaler and a visual granularity router, which are used to obtain multi-granularity visual features and select the appropriate visual granularity based on image and instruction, respectively. Besides, we introduce RGLF, which aligns router-predicted probabilities of multiple granularities with LMM preferences by a ranking loss, effectively helping the model learn to distinguish between different granularities based on image and instruction. Experimental results show that AVG-LLaVA not only exhibits superior performance across 11 benchmarks, but also significantly reduce the number of visual tokens and speed up inference in tasks that do not require fine-grained information. Particularly, on the AI2D benchmark, it reduces the number of visual tokens by 85.3% and speeds the inference by 2.53$\times$. We hope our work can inspire more attention to the visual granularity in LMMs.
While AVG-LLaVA has achieved good results, there is still considerable potential to be further explored. On text-intensive benchmarks, the model tends to select the finest-grained visual tokens, which may be due to the pooling directly reducing half of the tokens, resulting in significant differences in granularity size. In the future, we plan to design a more suitable granularity scaling network to provide richer visual granularities.
Appendix A Appendix
A.1 Training Details
We list the training hyperparameters for all stages in Table 5.
Hyperparameter Stage 1 Stage 2 Stage 3 Stage 4
Data size 595K 1M 1M 1M
Batch size 256 128 128 128
lr 1e-3 2e-5 1e-5 1e-3
Vision encoder lr - 2e-6 2e-5 -
@@ A.1 Training Details
DeepSpeed stage 2 3 3 3
Table 5: Hyperparameters of AVG-LLaVA.
A.2 Qualitative Evaluation

### v3

@@ 4.1 Settings
Table 1: Comparison with LMMs of the same size on general VQA benchmarks and text-oriented VQA benchmarks. The best results are marked in bold, and the second best results are underlined. Since MG-LLaVA is trained on significantly more data across two stages, we retrain it using the same data as ours for a fair comparison. We also explore the impact of additional two-stage training on the performance of LLaVA-NeXT using the same instruction fine-tuning data in Appendix A.4.
Training Datasets.
In the first training stage, since the real user interaction data used for visual instruction fine-tuning in LLaVA-NeXT are not open-sourced, we opt to extract 200K samples from the ALLaVA (Chen et al., 2024a) dataset as a substitute. Although LLaVA-NeXT replaces TextVQA (Singh et al., 2019) with DocVQA (Mathew et al., 2021) and SynDog-EN (Kim et al., 2022), the TextVQA has already been included in the training data of most existing LMMs. Consequently, we choose to retain it to ensure a fair comparison with other models. [^2] Our data recipe follows Open-LLaVA-NeXT (Lin and Long, 2024). In total, the visual instruction fine-tuning data we use contains 1M image-text pairs.
Implementation Details.
[…]
Evaluations.
We evaluate our model on three kinds of benchmarks: (1) general VQA benchmarks: GQA (Hudson and Manning, 2019), SciQA-Img (Lu et al., 2022), and VizWiz (Gurari et al., 2018); (2) text-oriented VQA benchmarks: TextVQA (Singh et al., 2019), ChartQA (Masry et al., 2022), DocVQA (Mathew et al., 2021), and AI2D (Kembhavi et al., 2016); and (3) general multimodal benchmarks: MME (Fu et al., 2023), MMB (Liu et al., 2023c), MMBCN (Liu et al., 2023c), POPE (Li et al., 2023c), and MMMU (Yue et al., 2023).
4.2 Main Results
@@
Table 2: Comparison with LMMs of the same size on general multimodal benchmarks.
General VQA Benchmarks.
The results in Table 1 show that AVG-LLaVA outperforms all standard-resolution LMMs on the general VQA benchmarks and achieves comparable performance to other high-resolution LMMs. Although it does not achieve the best results, it is important to note that AVG-LLaVA uses fewer visual tokens compared to other high-resolution models, and this comparison will be detailed in Section 4.3.
Text-oriented VQA Benchmarks.
In this category of benchmarks, as illustrated in Table 1, except for TextVQA, AVG-LLaVA outperforms all other comparison models. Back to TextVQA, AVG-LLaVA achieves the second-best performance, only trailing behind Mini-Gemini-HD. Notably, Mini-Gemini-HD utilizes more than twice the amount of data during the pretraining and approximately 1.5 times the amount of data during the visual instruction fine-tuning compared to AVG-LLaVA.
@@
[…]
General Multimodal Benchmarks.
Compared to traditional VQA datasets, this type of benchmarks cover a broader range of evaluation aspects, requiring models to possess more complex perception and reasoning capabilities. As summarized in Table 2, except for MME, AVG-LLaVA surpasses all other models across the remaining benchmarks, exhibiting superior overall performance and highlighting its adaptability and effectiveness across various disciplines. Specifically, AVG-LLaVA outperforms the second-best model by 6.1, 1.9, and 1.2 on MMEC, MMB, and MMBCN, respectively. Moreover, AVG-LLaVA’s performance on the POPE and MMMU benchmarks demonstrates its ability to reduce hallucinations and perform complex reasoning.
4.3 Computational Efficiency
To validate the effectiveness of dynamic visual granularity selection, we compare AVG-LLaVA with LLaVA-NeXT in terms of visual token number and inference speed across multiple benchmarks. Specifically, for each type of benchmarks, we select three benchmarks for comparison, and report the reduction in the number of visual tokens per grid and the actual speedup during inference.
As shown in Table 3, except for text-intensive VQA benchmarks that require very fine-grained visual information, such as TextVQA and ChartVQA, AVG-LLaVA significantly reduces the number of visual tokens and improves inference speed across other benchmarks. Particularly, on the AI2D benchmark, AVG-LLaVA achieves better performance than LLaVA-NeXT while using only 14.7% of visual tokens, and the inference speed increases by 2.53 $\times$. [^4] We also present qualitative results in Appendix A.5 and illustrate the effectiveness of adaptive visual granularity. Notably, even with the addition of two extra modules, there is no significant slowdown in inference speed on the ChartVQA benchmark when using a comparable number of visual tokens. Moreover, AVG-LLaVA only increases the number of parameters by 1.66% compared to LLaVA-NeXT.
@@ 4.4 Routing Visualization
To further understand the differences in the granularity selection of AVG-LLaVA across different benchmarks, we visualize the proportion of visual tokens selected at each granularity level for all benchmarks. Figure 4 shows the visualization results, it is evident that different tasks tend to favor different visual granularity, which is consistent with our expectations. In the case of text-intensive benchmarks like TextVQA, ChartQA, and DocVQA, the model requires fine-grained visual information, so the router predominantly selects the finest visual granularity. On the other hand, for benchmarks involving object-level questions, such as AI2D and MMMU, the model may find it easier to answer correctly by utilizing coarse-grained visual information. Although the 72 and 288-token granularities are seldom selected, their inclusion helps the model progressively learn and differentiate between various levels of visual granularity (see the ablation study in Section 4).
@@ 4.5 Ablation Study
Figure 5: Attention maps of AVG-LLaVA at different visual granularities. The instruction is “How many sheep are there? Answer the question with a single word.”. The number of visual tokens per grid selected by the router is 32.
Adaptive Visual Granularity vs. Fixed Visual Granularity.
@@
Router Granularity Selection vs. Random Granularity Selection.
@@
Impact of Router Input.
@@
Impact of Granularity Range.
@@
Impact of Router Training Methods.
We directly train the router using visual instructions fine-tuning with the cross-entropy loss function. Unlike our original approach where the router is directly supervised by the LMM feedback, this variant computes the loss on the LMM and backpropagates the gradient to the router using the Gumbel-Softmax technique (Jang et al., 2017). The results in Table 4(e) show that the LLM feedback allows the router to better distinguish the advantages and disadvantages of different granularities, thereby enabling it to select an appropriate granularity.
Importance of Ranking Granularity.
In Table 4(f) and Table 4(g), we remove the cross-entropy loss and ranking loss during the second stage, respectively. The results indicate that both types of loss are beneficial to model training and are complementary to each other, between which the ranking loss is more crucial. This underscores the necessity to train the router by ranking granularity based on LMM feedback.
4.6 Attention Map Visualization
To further understand how the appropriate granularity benefits the model in generating better answers, we visualize the attention map between the generated tokens and the visual tokens. The attention weights are calculated by accumulating the attention scores between image tokens and generated tokens across all layers and heads. As shown in Figure 5, when the instruction is “How many sheep are there? Answer the question with a single word.” the attention weights for the visual granularity selected by the router are mostly assigned to the two sheep, while the attention weights for other visual granularities are dispersed across the background. This means that selecting the appropriate visual granularity results in a more distinct attention map characterized by reduced background noise and enhanced focus on relevant regions, thereby improving model performance.
5 Conclusion
In this work, we propose AVG-LLaVA, an LMM that can adaptively select appropriate visual granularity based on input image and instruction. Besides, we introduce RGLF, which aligns router-predicted probabilities of multiple granularities with LMM preferences by a ranking loss, effectively helping the model learn to distinguish between different granularities. Experimental results show that AVG-LLaVA not only exhibits superior performance across 11 benchmarks, but also significantly reduce the number of visual tokens and speed up inference in tasks that do not require fine-grained information. In future work, we aim to develop different visual granularity scaling networks to obtain richer visual granularity and integrate the two-stage training into a single stage to improve efficiency.
Limitations
While AVG-LLaVA has achieved good results, there is still considerable potential to be further explored. On text-intensive benchmarks, the model tends to select the finest-grained visual tokens, which may be due to the pooling directly reducing half of the tokens, resulting in significant differences in granularity size. Designing a more suitable granularity scaling network to provide richer visual granularities may help alleviate this issue. Besides, the two-stage training introduces additional overhead, which could be alleviated by interleaving multi-granularity visual instruction fine-tuning and router training within a single stage.
Acknowledgments
The project was supported by National Key R&D Program of China (No. 2022ZD0160501), Natural Science Foundation of Fujian Province of China (No. 2024J011001), and the Public Technology Service Platform Project of Xiamen (No.3502Z20231043). We also thank the reviewers for their insightful comments.
Appendix A Appendix
A.1 Visual Granularity Selection Algorithm
In algorithm 1, we provide the detailed process of the router’s granularity selection.
A.2 Hyperparameter Analysis
@@
A.3 Training Details
We list the training hyperparameters for two stages in Table 5. Our setup mainly refers to LLaVA-NeXT Liu et al. (2024b).
Hyperparameter Stage 1 Stage 2
Data size 1M 1M
Batch size 128 128
lr 1e-5 1e-3
Vision encoder lr 2e-5 -
@@ A.3 Training Details
DeepSpeed stage 3 3
GPU 8 $\times$ H800 8 $\times$ H800
Training cost (#Hours) 65 14
Table 5: Training hyperparameters of AVG-LLaVA.
Model General VQA Text-oriented VQA
GQA ScienceQA VizWiz TextVQA ChartQA DocVQA AI2D
LLaVA-NeXT 64.2 70.1 57.6 64.9 54.8 74.4 66.6
LLaVA-NeXT-Extra 64.6 69.9 58.3 63.9 66.3 75.1 65.3
AVG-LLaVA 63.0 71.1 59.8 67.1 66.3 74.6 67.3
Table 6: Results on general VQA benchmarks and text-oriented VQA benchmarks. LLaVA-NeXT-Extra refers to training for two extra epochs on the same multimodal instruction-tuning data.
Model MME MMEC MMB MMBCN POPE MMMU
LLaVA-NeXT 1519.0 332.0 67.4 60.6 86.5 35.8
LLaVA-NeXT-Extra 1524.7 330.0 67.8 57.0 87.4 34.8
AVG-LLaVA 1557.4 366.8 69.9 61.8 87.4 37.4
Table 7: Results on general multimodal benchmarks.
A.4 Impact of Multiple Training Epochs on the Same Dataset
Since we conduct two additional training stages on the LLaVA-NeXT, we also investigate the impact of training LLaVA-NeXT for two extra epochs on the same multimodal instruction-tuning data.
From Tables 6 and 7, we can observe that although three additional two epoch trainings result in improvements on 7 benchmarks (e.g., ChartQA and DocVQA), there is a considerable performance decline on 6 benchmarks (e.g., TextVQA and MMBCN). This indicates that repeated training cannot improve the performance on all benchmarks. Besides, AVG-LLaVA performs better than LLaVA-NeXT-Extra on 9 benchmarks, is slightly worse on 2 benchmarks, and has a significant speed improvement, indicating that the advantage of AVG-LLaVA does not simply stem from repeated training.
A.5 Qualitative Evaluation
@@
A.6 Visualization of Router Selection for Different Instructions
Figure 11: Visualization of granularity selection using different instructions.
As shown in Figure 11, we input the same image with different instructions and then visualize the selected visual granularity on the image, i.e., the number of patches. As can be seen, even for the same image, the router selects different visual granularities for different instructions. When asking about the color of the car, the model does not require such fine-grained visual information, whereas when asking whether there is a cat, the model requires finer-grained visual information.